x^2-14x+49 is this problem factorable?

Answer by mmm4444bot(95)

Hello There:
x^2 - 14*x + 49 is factorable if we can find two numbers that multiply to make 49 and add to make -7.
With a little thought, you should realize that these two numbers are -7 and -7.
(-7)*(-7) = 49
-7 - 7 = -14
So, the factorization is:
(x - 7)*(x - 7) = (x - 7)^2
~ Mark